OpenStudy (anonymous):

In triangle ABC, line segment BC is extended through point C to point D. If m(angle)A=43 and m(angle)BCD=123, then m(angle)ABC= 1.)62 2.)80 3.)14 4.)166

OpenStudy (mathstudent55):

 If line segment BC is extended to point D, m<BCD = 180. How can it be 123?

OpenStudy (mathstudent55):

Is the angle that measures 123 BCD or ACD?

OpenStudy (anonymous):

Angle BCD=123

OpenStudy (mathstudent55):

Is the side that is extended BC or AC? In other words, did you copy the problem correctly?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Then this problem is incorrect. I think they want to make the angle that measures BCD an exterior angle. The measure of an exterior angle equals the sum of the measures of the remote interior angles. I think the intent of the problem may be to use the above fact about an exterior angle of a triangle. In this case this would be m<A + m<ABC = 123, and since m<A = 43, that makes m<ABC = 80. But there is an error with this problem as it is stated.

OpenStudy (anonymous):

This is the problem: In triangle ABC, line segment BC is extended through point C to point D. If m(angle)A=43 and m(angle)=123, then m(angle)ABC=

OpenStudy (anonymous):

the bc looks lilke this in thr problem" - BC

OpenStudy (mathstudent55):

BC with a bar on top? That just means line segment BC.

OpenStudy (anonymous):

yeah i know I dont get any of this!

OpenStudy (mathstudent55):

Wait, when you retyped the problem you left something out. Look a tthe last line: "If m(angle)A=43 and m(angle)=123, then m(angle)ABC=" What is m(angle)=123. Which angle has measure 123?

OpenStudy (anonymous):

angle BCD is 123

OpenStudy (anonymous):

angle A=43

OpenStudy (mathstudent55):

Right. That's the same you had originally. I still say there is an error in this problem.

OpenStudy (mathstudent55):

I think they meant m<ACD = 123. The m<BCD has to be 180, and that is only if Geometry allows for a 180-degree angle.

OpenStudy (anonymous):

okay so whats my answer?

OpenStudy (mathstudent55):

As the problem is stated there is no answer. If it's m<ACD = 123 instead of m<BCD = 123, then this is what you do: m<A + m<ABC = m<ACD 43 + m<ABC = 123 m<ABC = 123 - 43 m<ABC = 80

OpenStudy (anonymous):

<BCD=123 and <A=43... I need to find m<ABC

OpenStudy (mathstudent55):

Didn't I just do that just above? Of course, that is for the problem the way I modified it. The problem, as you wrote it, has no solution bec there is an error in it.

OpenStudy (anonymous):

okay thanks(:

OpenStudy (mathstudent55):

You're welcome. If you have a teacher for this, try to clear up this problem.